OpenStudy (anonymous):

What is the area enclosed by the curve 9x^2 + 25y^2 - 225 = 0?

OpenStudy (anonymous):

\[\left(\frac{x}{\frac{15}{3}}\right)^2 + \left(\frac{y}{\frac{15}{5}}\right)^2 = 1\] \[ \left(\frac{x}{5}\right)^2 + \left(\frac{y}{3}\right)^2 =1 \]